Listed below are pulse rates? (beats per? minute) from samples of adult males and females. Does there appear to be a? difference? Find the coefficient of variation for each of the two? samples; then compare the variation.

Male 90 71 64 72 72 53 65 52 84 69 64 60 96 54 62

Female 64 83 79 70 74 84 87 85 89 89 92 69 88 80 80

The coefficient of variation for the male pulse is ____ %

The coefficient of variation for the female pulse is ____ %

The coefficients of variation for each data set are ______________ Therefore, the male pulse rates vary ___________ the female pulse rates
Answer #1

For male \overline{x}=\frac{\sum xi}{n}=68.53 and s=\sqrt{\frac{\sum (xi-\overline{x})^2}{n-1}}=13.0377

Hence coefficient of variation for male is \frac{s}{\overline{x}}=\frac{13.0377}{68.5333}=0.1902=19.02\%

Similarly for female \overline{y}=\frac{\sum yi}{n}=80.8667 and s=\sqrt{\frac{\sum (yi-\overline{y})^2}{n-1}}=8.3312

Hence coefficient of variation for male is \frac{s}{\overline{y}}=\frac{8.3312}{80.8667}=0.1030=10.30\%

The coefficients of variation for each data set are different. Therefore, the male pulse rates vary significantly from the female pulse rates

And